what do you do when your are trying to find a missig measurement for an angle what is the formula for it for example 50 degreese , 40 degreese and __ degreese what do you do to find the measure of the angle
Answer by benni1013(206)

The formula is a+b+c=180°. So 50+40+c=180°==>90+c=180==>180-90=c==>90=c